### v2

@@ 5 Inexact proximal point method (iPPM) and its convergence
To prove Theorem 5.3, we first introduce an important inequality characterizing the one-step improvement, which was first established in [33, Equation 2.7]. We provide some details in Section D.2 for completeness.
Lemma 5.1 (Inexact one-step improvement [33, Equation 2.7]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by the iPPM 20 under B with the positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then there exists a $\hat{k}\geq 0$ such that
@@
$\|x_{k}-x^{\star}\|<\frac{r}{2},$ $\;\forall k\geq k_{1},\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|<\frac{r}{2},\;\forall k\geq k_{2}.$
@@
which implies the point $\text{prox}_{c_{k}f}(x_{k})$ satisfies 23. The same analysis in Theorem 4.2 shows
@@
Consequently, combining 24 with 25 yields
@@
Finally, as we know $\delta_{k}\to 0$ and $\theta_{k}<1,\;\forall k\geq\max\{k_{1},k_{2}\},$ there exists a $\bar{k}\geq\max\{k_{1},k_{2},k_{3}\}$ such that
@@
In Theorem 5.3, criterion A serves to guarantee that the iterates can reach the neighborhood $\mathcal{U}$ in 21 (cf. the asymptotic convergence from Theorem 5.1). If 21 holds globally (i.e., $r=+\infty$), the same linear convergence result holds with B only. Note that the convergence proof in Theorem 5.3 is very modular. Indeed, thanks to the regularity conditions (i.e., 21), our proof appears simpler and less conservative than typical proofs in the literature; see the discussions after Theorem 4.2.
@@ 6 Applications
In this section, we present numerical experiments demonstrating the linear convergence of PPM. Throughout the experiment, we use the modeling package YALMIP [48] to formulate the subproblem 12 and call the conic solver MOSEK [49] to solve it. Our code is available at
@@
We consider three different applications of convex optimization in machine learning and signal processing: linear support vector machine (SVM) ([50]), lasso ([51]), and elastic-net ([52]) respectively. The problem formulations are
@@
For linear SVM, we consider three datasets (“a1a”, “Australian”, “fourclass”) from [53]. The regularization term is set as $\rho=1.$ The step size parameter of the PPM is chosen as $c_{k}=1$ for all $k\geq 1$. For Lasso, we randomly generate $A\in\mathbb{R}^{n\times m}$ and $\hat{x}\in\mathbb{R}^{m}$ with $s$ elements being zero ($s<m$), then construct the vector $y=A\hat{x}$. We consider problems with three different sizes: $(n=10,m=40,s=5),(n=20,m=50,s=10),$ and $(n=30,m=60,s=15)$. The regularization terms are set as $\rho=10$. The step size parameter of the PPM is chosen as $c_{k}=0.16$ for all $k\geq 1$. For Elastic-Net, we consider the same data in Lasso with the regularization terms $\rho=10$ and $\mu=1.$ These problems satisfy the regularity conditions in Theorem 4.2. For each application, we run the PPM on three different data sets. The numerical results are shown in Figure 3, which confirms the fast linear convergence of the PPM.
@@ 7 Conclusion
We have also presented simple and modular proofs for the exact PPM for both convex and weakly convex functions which makes the analysis of the exact PPM more accessible (see Theorems 4.2 and 4.3). In the setting of inexact PPM, we first clarified the subtly in terms of inexactness and infeasibility, and then established the linear convergence of inexact PPM when the inexactness is controlled properly (see Theorem 5.3). The numerical experiments also validate our theoretical findings. We believe these results on (inexact) PPM will facilitate algorithm development in nonsmooth optimization, and we are particularly interested in further applications in large-scale conic optimization [54, 55].
@@ Appendix A Subdifferential characterization of weakly convex functions
We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is proper if $f(x)<\infty$ for at least one $x\in\mathbb{R}^{n}$ and $-\infty<f(x)$ for all $x\in\mathbb{R}^{n}.$ We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is closed if the epigraph $\text{epi}f:\{(x,\alpha)\in\mathbb{R}^{n}\times\mathbb{R}\mid f(x)\leq\alpha\}$ is closed.
Definition A.1.
For a closed function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$, we define the Fréchet subdifferential as
^$\hat{\partial}f(x)=\mathopen{}\mathclose{{}\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{}\left\langle s,y-x}\right\rangle}{\|y-x\|}\geq 0}\right\}.$
Definition A.2.
@@
For this class of $\rho$-weakly convex functions, the Fréchet subdifferential always exists for any point $x\in\mathbb{R}^{n}$. This is because the Fréchet subdifferential is the same as the Clarke subdifferential for $\rho$-weakly convex functions and the Clarke subdifferential always exists [37, Section 3 and fact 5]. The class of $\rho$-weakly convex functions also has the following nice characterizations. These characterizations offer favorable properties for $\rho$-weakly convex functions, which can be viewed as extensions from strongly convex functions.
Lemma A.1 (Subdifferential characterization [26, Lemma 2.1]).
The following statements are equivalent for any closed function $f$.
@@
    $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y)+$ $\frac{\rho\lambda(1-\lambda)}{2}\|x-y\|^{2},$
    $\forall x,y\in\mathbb{R}^{n},\lambda\in[0,1].$
@@
    $\mathopen{}\mathclose{{}\left\langle v-w,x-y}\right\rangle\geq-\rho\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{n},v\in\hat{\partial}f(x),\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ w\in\hat{\partial}f(y).$
@@
The subgradient inequality 26 is particularly useful in our proof for Theorem 3.1. The proof repeatedly applies 26 to establish the connection between subdifferential and cost value gap.
@@ Appendix B Proofs in Section 2
Assume that RSI 2 holds with constant $\mu_{\mathrm{r}}>0$. We first note that $\mu_{\mathrm{r}}\leq L$. Indeed, this can be seen from the following fact: let $x\in\mathbb{R}^{n},x\notin S,\hat{x}\in S$, we have
@@
where the first inequality uses RSI 2, and the last inequality applies Cauchy-Schwarz and 27.
@@
where 29a plugs in definition of $x_{k+1}$ and uses the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\hat{x}_{k}\|$, 29b simply expends the square, and 29c uses 2 to upper bound the inner product term and applies the $L$-smoothness inequality 27 to $\|\nabla f(x_{k})\|=\|\nabla f(x_{k})-\nabla f(\hat{x}_{k})\|$ $\leq L\|x_{k}-\hat{x}_{k}\|$ (noting that $\nabla f(\hat{x}_{k})=0$).
@@
where the first inequality applies 28. If the step size $t_{k}$ is chosen such that $0<t_{k}<\frac{2}{L}$, then $\frac{-2t_{k}+Lt_{k}^{2}}{2}<0$. Then, applying PL 3 to the above inequality and subtracting $f^{\star}$ from both sides, we arrive at the linear convergence in the cost value gap
@@
To summarize, if we choose the step size $t_{k}$ such that the constants in 29d and 30 less than one (equivalently, $10\leq 1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}<$ 1 and $0<t_{k}<\frac{2}{L}$), linear convergence for both the distance and the cost value gap is guaranteed. In particular, choosing $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$, we have $11-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=1-\frac{\mu_{\mathrm{r}}}{L}<$ 1 and $t_{k}\leq\frac{L}{L^{2}}<\frac{2}{L}$ as $\mu_{\mathrm{r}}\leq L$, leading to the desired linear convergence:
$\mathrm{dist}^{2}(x_{k+1},S)$ $\leq(1-\mu_{\mathrm{r}}/L)\mathrm{dist}^{2}(x_{k},S),$
@@
Note that if 3 holds with $\mu_{\mathrm{p}}>0$, then $\mu_{\mathrm{p}}\leq\frac{L^{3}}{2\mu_{\mathrm{r}}L-\mu_{\mathrm{r}}^{2}}$. Otherwise, the above cost value decrease would contradict with the fact that the cost value gap is always nonnegative. This choice of step size $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$ is not new, which was used in [34, Proposition 1] or implicitly [35, Proposition 1].
Remark B.1.
Note that Theorem 2.1 is consistent with the linear convergence result in [10, Theorem 1] since the step size $t_{k}=\frac{1}{L}$ also satisfies $0<t_{k}<\frac{2}{L}$. Despite choosing $t_{k}=\frac{1}{L}$ renders a cleaner reduction constant $(1-\frac{\mu_{\mathrm{r}}}{L})$ in [10, Theorem 1], it is unclear if the distance also converges linearly under this stepsize choice (i.e., it is not guaranteed that $t_{k}=\frac{1}{L}$ leads to the constant $11-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=2-2\frac{\mu_{\mathrm{r}}}{L}<$ 1).
@@
We note that the following inequality holds for any $\bar{x}\in\mathbb{R}^{n}$ with a finite function value $f(\bar{x})$ [21, Section 2]
@@
$\overline{|\nabla f|}(\bar{x}):=\liminf_{x\to\bar{x},f(x)\to f(\bar{x})}|\nabla f|(x).$ (32)
The result from [38, Proposition 8.5] has shown that
@@
where $\partial f(\bar{x})$ is the limiting subdifferential [37, Equation 18] at $\bar{x}$ defined as
$\partial f(\bar{x})=\{s\in\mathbb{R}^{n}\mid\exists x_{k}\to\bar{x}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ s_{k}\in\hat{\partial}f(x_{k})\leavevmode\nobreak\ \text{such that}\leavevmode\nobreak\ s_{k}\to s\}.$
Note that we use $\partial f$ to denote both the convex subdifferential and limiting subdifferential of $f$. Then we claim that for the class of functions whose Fréchet subdifferential and limiting subdifferential coincide, we have $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x})).$ Indeed, suppose ^$\hat{\partial}f(\bar{x})=\partial f(\bar{x})$, using 33 and 31, we have
@@
On the other hand, $\overline{|\nabla f|}(\bar{x})\leq|\nabla f|(\bar{x})$ always holds true by defintion 32. Thus, $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$. One sufficient condition for Fréchet subdifferential and limiting subdifferential being the same (i.e., ^$\hat{\partial}f(\bar{x})=\partial f(\bar{x})$) is $f$ being subdifferentially regular at $\bar{x}$ [37, Definition 2 and fact 5]. For smooth, convex, and weakly convex functions, they are subdifferentially regular at every point [37, Page 24]. Thus, the result follows.
@@
To finish the proof for 15b with coefficient $\theta_{k}\leq 1/{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}$. We first review a crucial but standard result about the proximal mapping.
Lemma C.1 ([16, Equation 3.1]).
For any $c_{k}>0$ in 12, the following holds true
@@
As shown in [56], the proximal point operator for a convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ with $\alpha>0$ is firmly nonexpensive, i.e., $\forall x,y\in\mathbb{R}^{n},$
@@
where the last inequality comes from the fact that the inner product is nonnegative by 35.
Setting $x=x_{k}$ and $y=\Pi_{S}(x_{k})$, we have $x_{k+1}=P(x_{k})$, and $y=P(y)$, which leads to the desired result in 34. ∎
@@
where the second inequality is from 34, the third inequality comes from the optimality condition 16, and the last inequality uses EB. Rearranging terms, we have

### v3

@@ 5 Inexact proximal point method (iPPM) and its convergence
To prove Theorem 5.3, we first introduce an important inequality characterizing the one-step improvement, which was first established in [luque1984asymptotic, Equation 2.7]. We provide some details in Section D.2 for completeness.
Lemma 5.1 (Inexact one-step improvement [luque1984asymptotic, Equation 2.7]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by the iPPM 23 under B with the positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then there exists a $\hat{k}\geq 0$ such that
@@
$\|x_{k}-x^{\star}\|<\frac{r}{2},$ $\;\forall k\geq k_{1},\penalty 10000\ \text{and}\penalty 10000\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|<\frac{r}{2},\;\forall k\geq k_{2}.$
@@
which implies the point $\text{prox}_{c_{k}f}(x_{k})$ satisfies 26. The same analysis in Theorem 4.2 shows
@@
Consequently, combining 27 with 28 yields
@@
Finally, as we know $\delta_{k}\to 0$ and $\theta_{k}<1,$ for all $k\geq\max\{k_{1},k_{2}\},$ there exists a $\bar{k}\geq\max\{k_{1},k_{2},k_{3}\}$ such that
@@
In Theorem 5.3, criterion A serves to guarantee that the iterates can reach the neighborhood $\mathcal{U}$ in 24 (cf. the asymptotic convergence from Theorem 5.1). If 24 holds globally (i.e., $r=+\infty$), the same linear convergence result holds with B only. Note that the convergence proof in Theorem 5.3 is very modular. Indeed, thanks to the regularity conditions (i.e., 24), our proof appears simpler and less conservative than typical proofs in the literature; see the discussions after Theorem 4.2.
@@ 6 Applications
In this section, we present numerical experiments demonstrating the linear convergence of PPM. Throughout the experiment, we use the modeling package YALMIP [Lofberg2004] to formulate the subproblem 15 and call the conic solver MOSEK [mosek] to solve it. Our code is available at
@@
We consider three different applications of convex optimization in machine learning and signal processing: linear support vector machine (SVM) [zhang2015stochastic], lasso [tibshirani1996regression], and elastic-net [zou2005regularization], respectively. The problem formulations are
@@
For linear SVM, we consider three datasets (“a1a”, “Australian”, “fourclass”) from [chang2011libsvm]. The regularization term is set as $\rho=1.$ The step size parameter of the PPM is chosen as $c_{k}=1$ for all $k\geq 1$. For Lasso, we randomly generate $A\in\mathbb{R}^{n\times m}$ and $\hat{x}\in\mathbb{R}^{m}$ with $s$ elements being zero ($s<m$), then construct the vector $y=A\hat{x}$. We consider problems with three different sizes: $(n=10,m=40,s=5),(n=20,m=50,s=10),$ and $(n=30,m=60,s=15)$. The regularization terms are set as $\rho=10$. The step size parameter of the PPM is chosen as $c_{k}=0.16$ for all $k\geq 1$. For Elastic-Net, we consider the same data in Lasso with the regularization terms $\rho=10$ and $\mu=1.$ These problems satisfy the regularity conditions in Theorem 4.2. For each application, we run the PPM on three different data sets. The numerical results are shown in Figure 3, which confirms the fast linear convergence of the PPM.
@@ 7 Conclusion
We have also presented simple and modular proofs for the exact PPM for both convex and weakly convex functions which makes the analysis of the exact PPM more accessible (see Theorems 4.2 and 4.3). In the setting of inexact PPM, we first clarified the subtly in terms of inexactness and infeasibility, and then established the linear convergence of inexact PPM when the inexactness is controlled properly for convex optimization (see Theorem 5.3). The numerical experiments also validate our theoretical findings. We believe these results on (inexact) PPM will facilitate algorithm development in nonsmooth optimization, and we are particularly interested in further applications in large-scale conic optimization [zheng2021chordal, zheng2020chordal].
@@ Appendix A Subdifferential characterization of weakly convex functions
We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is proper if $f(x)<\infty$ for at least one $x\in\mathbb{R}^{n}$ and $\infty<f(x)$ for all $x\in\mathbb{R}^{n}.$ We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is closed if the epigraph $\text{epi}f:\{(x,\alpha)\in\mathbb{R}^{n}\times\mathbb{R}\mid f(x)\leq\alpha\}$ is closed.
Definition 2.
For a closed function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $x$ such that $f(x)$ is finite, we define the Fréchet subdifferential at the point $x$ as
$\hat{\partial}f(x)=\mathopen{}\mathclose{{\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle}{\|y-x\|}\geq 0}}\right\}.$
Definition 3.
@@
For this class of $\rho$-weakly convex functions, the Fréchet subdifferential always exists for any point $x\in\mathbb{R}^{n}$. This is because the Fréchet subdifferential is the same as the Clarke subdifferential for $\rho$-weakly convex functions and the Clarke subdifferential always exists [li2020understanding, Section 3 and fact 5]. The class of $\rho$-weakly convex functions also has the following nice characterizations. These characterizations offer favorable properties for $\rho$-weakly convex functions, which can be viewed as extensions from strongly convex functions.
Lemma A.1 (Subdifferential characterization [davis2019stochastic, Lemma 2.1]).
The following statements are equivalent for any closed and proper function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$.
@@
    $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y)+$ $\frac{\rho\lambda(1-\lambda)}{2}\|x-y\|^{2},\;\forall x,y\in\mathbb{R}^{n},\lambda\in[0,1].$
@@
    $\mathopen{}\mathclose{{\left\langle v-w,x-y}}\right\rangle\geq-\rho\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{n},v\in\hat{\partial}f(x),\penalty 10000\ \text{and}\penalty 10000\ w\in\hat{\partial}f(y).$
@@
The subgradient inequality 29 is particularly useful in our proof for Theorem 3.1. The proof repeatedly applies 29 to establish the connection between subdifferential and cost value gap.
@@ Appendix B Proofs in Section 2
Assume that RSI 2 holds with constant $\mu_{\mathrm{r}}>0$. We first note that $\mu_{\mathrm{r}}\leq L$. Indeed, this can be seen from the following fact: let $x\in\mathbb{R}^{n},x\notin S,\hat{x}\in\Pi_{S}(x)$, we have
@@
where the first inequality uses RSI 2, and the last inequality applies Cauchy-Schwarz and 30.
@@
where 32a plugs in definition of $x_{k+1}$ and uses the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\hat{x}_{k}\|$, 32b simply expends the square, and 32c uses 2 to upper bound the inner product term and applies the $L$-smoothness inequality 30 to $\|\nabla f(x_{k})\|=\|\nabla f(x_{k})-\nabla f(\hat{x}_{k})\|$ $\leq L\|x_{k}-\hat{x}_{k}\|$ (noting that $\nabla f(\hat{x}_{k})=0$).
@@
where the first inequality applies 31. If the step size $t_{k}$ is chosen such that $0<t_{k}<\frac{2}{L}$, then $\frac{-2t_{k}+Lt_{k}^{2}}{2}<0$. Then, applying PL 3 to the above inequality and subtracting $f^{\star}$ from both sides, we arrive at the linear convergence in the cost value gap
@@
To summarize, if we choose the step size $t_{k}$ such that the constants in 32d and 33 less than one (equivalently, $0\leq 1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}<1$ and $0<t_{k}<\frac{2}{L}$), linear convergence for both the distance and the cost value gap is guaranteed. In particular, choosing $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$, we have $1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=1-\frac{\mu_{\mathrm{r}}^{2}}{L^{2}}<1$ and $t_{k}\leq\frac{L}{L^{2}}<\frac{2}{L}$ as $\mu_{\mathrm{r}}\leq L$, leading to the desired linear convergence:
$\mathrm{dist}^{2}(x_{k+1},S)$ $\leq(1-\mu_{\mathrm{r}}^{2}/L^{2})\mathrm{dist}^{2}(x_{k},S),$
@@
Note that if 3 holds with $\mu_{\mathrm{p}}>0$, then $\mu_{\mathrm{p}}\leq\frac{L^{3}}{2\mu_{\mathrm{r}}L-\mu_{\mathrm{r}}^{2}}$. Otherwise, the above cost value decrease would contradict the fact that the cost value gap is always nonnegative. This choice of step size $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$ is not new, which was used in [guille2022gradient, Proposition 1] or implicitly [zhang2020new, Proposition 1].
Remark 2.
Note that Theorem 2.1 is consistent with the linear convergence result in [karimi2016linear, Theorem 1] since the step size $t_{k}=\frac{1}{L}$ also satisfies $0<t_{k}<\frac{2}{L}$. Despite choosing $t_{k}=\frac{1}{L}$ renders a cleaner reduction constant $(1-\frac{\mu_{\mathrm{r}}}{L})$ in [karimi2016linear, Theorem 1], it is unclear if the distance also converges linearly under this stepsize choice (i.e., it is not guaranteed that $t_{k}=\frac{1}{L}$ leads to the constant $1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=2-2\frac{\mu_{\mathrm{r}}}{L}<1$).
@@
We note that the following inequality holds for any $\bar{x}\in\mathbb{R}^{n}$ with a finite function value $f(\bar{x})$ [drusvyatskiy2021nonsmooth, Section 2]
@@
$\overline{|\nabla f|}(\bar{x})=\liminf_{x\to\bar{x},f(x)\to f(\bar{x})}|\nabla f|(x).$ (35)
The result from [ioffe2017variational, Proposition 8.5] has shown that
@@
where $\partial f(\bar{x})$ is the limiting subdifferential [li2020understanding, Equation 18] at $\bar{x}$ defined as
$\partial f(\bar{x})=\{s\in\mathbb{R}^{n}\mid\exists x_{k}\to\bar{x}\penalty 10000\ \text{and}\penalty 10000\ s_{k}\in\hat{\partial}f(x_{k})\penalty 10000\ \text{such that}\penalty 10000\ s_{k}\to s\}.$
Then we claim that for the class of functions whose Fréchet subdifferential and limiting subdifferential coincide, we have $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x})).$ Indeed, suppose $\hat{\partial}f(\bar{x})=\partial f(\bar{x})$, using 36 and 34, we have
@@
On the other hand, $\overline{|\nabla f|}(\bar{x})\leq|\nabla f|(\bar{x})$ always holds true by defintion 35. Thus, $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$.
Below, we show that Fréchet subdifferential and limiting subdifferential match for 1) closed and smooth; 2) closed and convex; 3) closed and weakly convex functions. We also assume that the function is proper. We first show that the result holds for a $\rho$-weakly convex closed function $f:\mathbb{R}\to\overline{\mathbb{R}}$, and the result for a proper convex closed function follows by setting $\rho=0$. By the definition of $\partial f(\bar{x})$, it is clear that $\hat{\partial}f(\bar{x})\subseteq\partial f(\bar{x})$ (choosing $x_{k}=\bar{x}$ and $s_{k}=s\in\hat{\partial}f(\bar{x})$). On the other hand, if $s\in\partial f(\bar{x})$, then we can find $\{x_{k}\}\to\bar{x},\{s_{k}\}\to s,$ and $s_{k}\in\hat{\partial}f(x_{k})$. Thus, for all $y\in\mathbb{R}^{n}$, we have
$\mathopen{}\mathclose{{\left\langle s,y-\bar{x}}}\right\rangle-\frac{\rho}{2}\|y-\bar{x}\|^{2}$ $=\lim_{k\rightarrow+\infty}\mathopen{}\mathclose{{\left\langle s_{k},y-x_{k}}}\right\rangle-\frac{\rho}{2}\|y-x_{k}\|^{2}$
$=\liminf_{k\rightarrow+\infty}\mathopen{}\mathclose{{\left\langle s_{k},y-x_{k}}}\right\rangle-\frac{\rho}{2}\|y-x_{k}\|^{2}$
$\overset{(a)}{\leq}\liminf_{k\rightarrow+\infty}f(y)-f(x_{k})$
$\overset{(b)}{\leq}f(y)-f(\bar{x}),$
where $(a)$ applies the subgradient inequality 29, and $(b)$ uses the closeseness of $f$ (i.e., lower semicontinuity $\liminf_{y\to\bar{x}}f(y)\geq f(\bar{x})$). Thus, Fréchet subdifferential and limiting subdifferential match for $\rho$-weakly convex closed functions.
For a smooth function $f$, it is also clear that $\{\nabla f(\bar{x})\}\subseteq\partial f(\bar{x})$ since $\{\nabla f(\bar{x})\}\subseteq\hat{\partial}f(\bar{x})\subseteq\partial f(\bar{x})$. On the other hand, let $s\in\partial f(\bar{x})$. We have $x_{k}\to\bar{x}$ and $\{\nabla f(x_{k})\}\to s.$ Since $f$ is smooth, the gradient $\nabla f$ is continuous, implying $\{\nabla f(x_{k})\}\to\nabla f(\bar{x})$. Thus, we have $s=\nabla f(x)$. This proves that $\{\nabla f(\bar{x})\}=\hat{\partial}f(\bar{x})=\partial f(\bar{x})$. Thus, the result follows.
@@
To finish the proof for 18b with coefficient $\theta_{k}\leq 1/{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}$. We first review a crucial but standard result about the proximal mapping.
Lemma C.1 ([leventhal2009metric, Equation 3.1]).
For any $c_{k}>0$ in 15, the following holds true
@@
As shown in [rockafellar2021characterizing], the proximal point operator for a convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ with $\alpha>0$ is firmly nonexpensive, i.e., $\forall x,y\in\mathbb{R}^{n},$
@@
where the last inequality comes from the fact that the inner product is nonnegative by 38.
Setting $x=x_{k}$ and $y=\Pi_{S}(x_{k})$, we have $x_{k+1}=P(x_{k})$, and $y=P(y)$, which leads to the desired result in 37. ∎
@@
where the second inequality is from 37, the third inequality comes from the optimality condition 19, and the last inequality uses EB. Rearranging terms, we have